OpenStudy (anonymous):

Solve the following system of equations. 8x-7y=-21 -3x-5y=-15 x=? y=?

OpenStudy (anonymous):

make in your equation appear 3x ...how will you do that?

OpenStudy (anonymous):

first*

OpenStudy (anonymous):

3x(8x-7y=-21) ?

OpenStudy (anonymous):

you mean multiply both sides by 3x?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

then you will have 24x...you want that?

OpenStudy (anonymous):

i dont know!

OpenStudy (anonymous):

if i knew the answer, i wouldnt be asking the question..

OpenStudy (anonymous):

why would I want 3x to appear in the 1st equation?

OpenStudy (anonymous):

if you find that ..you will understand whole problem

OpenStudy (anonymous):

you told me to put 3x in the equation so thats what i did.. why is that wrong now?

OpenStudy (anonymous):

no..not put 3x in the 1st equation...you need to make 3x appear in the 1st equation...what is the little trick that will make the 1st equation have 3x instead of 8x?

OpenStudy (anonymous):

wouldnt it be easier to just make both equations have 24x?

OpenStudy (anonymous):

thats okay and it is correct

OpenStudy (anonymous):

so how can u make both equations have 24x?

OpenStudy (anonymous):

multipy the top by 3 and the bottom by -8. 24x-21y=-61 and 24x+40y=120

OpenStudy (anonymous):

one small mistake there 21*3=63 not 61

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now you have 24x in both of your equations

OpenStudy (anonymous):

yes. now what

OpenStudy (anonymous):

solve the 1st equation for 24x and then put it in your 2nd equation

OpenStudy (anonymous):

idk how. i moved 21y over. and made it 24x=-63+21y and then tried to divide by 24 but it doesnt work out

OpenStudy (anonymous):

no you dont need to divide this ...you need this!

OpenStudy (anonymous):

put the right hand part of this equation on your 2nd equation

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

your 2nd equation is 24x+40y=120 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

in that 2nd equation where you see 24x put what you found before from the first

OpenStudy (anonymous):

i didnt find anything from the first..?

OpenStudy (anonymous):

all i have from the first is 24x=-63+21y

OpenStudy (anonymous):

what you need is to get rid of x so you can have an equation with only y as unknown

OpenStudy (anonymous):

you have 24x from the first:)

OpenStudy (anonymous):

doesnt matter now. my page with the question timed out.

OpenStudy (anonymous):

what do you mean?you dont need to solve that anymore?

OpenStudy (anonymous):

nope. we took to long.

OpenStudy (anonymous):

look...from the first equation you have -63+21y=24x from the 2nd you have 24x+40y=120...so the 2nd equation becomes -63+21y+40y=120 solve that to find y

OpenStudy (anonymous):

61y=183--->y=183/61 calculator and get your result

OpenStudy (anonymous):

y=3